OpenStudy (naate):

x^2+b/ax+(b/2a)^2=-c/a....i need help :D #Medal Dani_rose

OpenStudy (naate):

@Dani_rose

OpenStudy (dani_rose):

\[x^2+\frac{ b }{ ax }+\frac{ b }{ 2a }^2=\frac{ -c }{ a }\]

OpenStudy (naate):

yea, but the b/2a ^2 is in parenthesis

OpenStudy (naate):

not the ^2 though, sorry

OpenStudy (dani_rose):

ok

OpenStudy (naate):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

can you screenshot the question

OpenStudy (naate):

not on my compter.....i dont think @rebeccaxhawaii